OpenStudy (anonymous):

Find the value of the unknown side length for the given perimeter P. P = 28 m A. b = 40 m B. b = 22 m C. b = 6 m D. b = 5 m

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

perimeter = sum of all sides of the figure so basically you add up all the sides to get the perimeter

jimthompson5910 (jim_thompson5910):

b+6+11+5 = 28 solve for b

OpenStudy (anonymous):

So the answer is b.?

jimthompson5910 (jim_thompson5910):

you got 22?

jimthompson5910 (jim_thompson5910):

b+6+11+5 = 28 b+17+5 = 28 b+22 = 28 b = ???

OpenStudy (anonymous):

22 i got it

jimthompson5910 (jim_thompson5910):

b = 22 is incorrect